### v1

@@ Abstract
Bimanual manipulation tasks typically involve multiple stages which require efficient interactions between two arms, posing step-wise and stage-wise challenges for imitation learning systems. Specifically, failure and delay of one step will broadcast through time, hinder success and efficiency of each sub-stage task, and thereby overall task performance. Although recent works have made strides in addressing certain challenges, few approaches explicitly consider the multi-stage nature of bimanual tasks while simultaneously emphasizing the importance of inference speed. In this paper, we introduce a novel keypose-conditioned consistency policy tailored for bimanual manipulation. It is a hierarchical imitation learning framework that consists of a high-level keypose predictor and a low-level trajectory generator. The predicted keyposes provide guidance for trajectory generation and also mark the completion of one sub-stage task. The trajectory generator is designed as a consistency model trained from scratch without distillation, which generates action sequences conditioning on current observations and predicted keyposes with fast inference speed. Simulated and real-world experimental results demonstrate that the proposed approach surpasses baseline methods in terms of success rate and operational efficiency.
@@ 1 Introduction
Bimanual manipulation are ubiquitous in daily activities and industrial applications such as cloth folding, battery slotting and tool assembling [2, 33]. These tasks typically require intricate coordination between two arms and involve multiple stages, resulting in more complexity than unimanual case  [7]. One intuitive solution is to assign a specific role to each arm based on human-specified rules, with each arm being programmed independently [17, 10]. However, this approach proves to be labor-intensive and hard-coded, as coordination rules often require customization for specific tasks, particularly those involving multiple stages [19]. Recently, imitation learning (IL), a.k.a. learning from demonstration (LfD), has been widely applied in bimanual manipulation. IL is a straightforward yet effective approach that obtains automated agents in an end-to-end manner by imitating demonstrations, thereby eliminating the need for explicitly and manually designing coordination rules [3, 32, 18].
@@ 2 Related Work
IL enables robots to acquire skills by learning directly from experts’ demonstrations. The intuitive process makes it widely used in robotic manipulation. IL in manipulation can be categorized into non-parametric approaches (e.g., Dynamic Movement Primitives [12]) and parametric methods promising for high-dimensional tasks [32, 29]. Behavioral cloning (BC) [4] is a classic parametric IL algorithm that acts similarly as supervised learning, directly fitting the mapping from observations (inputs) to actions (labels). Despite its simplicity and efficacy, BC-based approaches are susceptible to compounding errors [33] and multi-modality in demonstrations [1], which is worse when it comes to bimanual cases. Recently, the integration of generative modeling techniques into IL has brought success [2, 33, 1, 25]. Predicting action sequences in these works mitigates compounding errors by shortening the effective horizon [33, 1, 25]. Some works [33, 25] address the multi-modality challenge by leveraging Variational Auto-Encoder (VAE) [16] to focus on the predominant behavior style. Conversely, other works capture multi-modal action distributions by modeling the policy as a diffusion model [1, 11, 28]. However, diffusion models inherently incur significant inference latency due to iterative denoising, which presents a critical concern especially in real-world deployment. In this work, we focus on a variant of diffusion models, Consistency Models (CMs) [27, 26], to formulate robot policies. CMs can capture multi-modal demonstrations like DP [27] while generating action sequences with only one-step inference, i.e., fast inference speed. Some concurrent works [21, 24] also apply CM to robotic manipulation tasks with the aim of accelerating run-time inference. However, these approaches distill CMs from well-trained diffusion models (i.e., consistency distillation (CD) [27]) and focus only on unimanual tasks. Here, CD necessitates additional efforts to learn a diffusion model first and depends on the quality of the learned diffusion model. To this end, we adopt consistency training (CT) [27, 26] in this paper. Through comprehensive experiments, we demonstrate that CMs trained from scratch can achieve high-performance deployment and real-time reaction in bimanual tasks.
Hierarchical IL for Multi-stage Tasks.
@@
Hierarchical IL (HIL) is proposed to enhance the performance for multi-stage tasks, by decomposing a complex task into smaller and more manageable ones. One line of studies pre-define a set of manipulation primitives, such as picking, placing, moving and stabilizing. Then a high-level policy is learned to sequentially choose primitives to be executed [22]. In bimanual cases, this intuition is implemented by assigning a fixed role for each arm, after which each arm learns an independent policy for its designated role [10]. In this way, the bimanual system is decomposed into two unimanual cases. However, such a method is constrained to specific types of tasks. Another line of work segments the demonstration into sub-stage trajectories. For instance, [34] employs a bottom-up clustering on observations to accomplish unsupervised segmentation, while [23] establishes heuristic rules based on robot kinematics. One technique to segment trajectories is extracting keyposes. Keypose typically refers to the next-best end-effector (EE) pose or joint positions where the arm reaches a specific state, e.g., EE velocity approaching zero or the gripper beginning to open or close. With identified keyposes, IL on a multi-stage task comes down to modeling both the high-level keypose transition and the low-level sub-trajectories connecting two-keypose pairs. Built on this idea, several works have made achievements in unimanual manipulation [9, 31, 23]. However, applying keypose-based approaches in bimanual tasks poses a challenge in coordinating two arms. In this study, we address this issue by adopting a merging method. Furthermore, we condition the learning of the trajectory generator on target keyposes, to improve per-stage reliability and enhance overall success rates.
@@ 3 Preliminaries: Consistency Models as Visuomotor Policy
Consistency Models in Generative Modeling.
CMs are proposed to remove the iterative sampling process in DDPMs to accelerate generation speed [27] as another score-based generative modeling method. The core of CMs is the existence of an ordinary differential equation (ODE) shares the same per-step distribution $p_{\sigma}({\mathbf{x}})$ of samples ${\mathbf{x}}^{\sigma}$ with diffusion SDE [28]:
@@
where ${\bm{f}}(\cdot,\cdot)$ is the consistency function mapping diffused samples to its origin.
Training and Sampling of Consistency Models.
Training CMs is to approximate the consistency function ${\bm{f}}$ with a parameterized one ${\bm{f}}_{\bm{\theta}}$ (typically neural networks) by learning to enforce eq. 2. Recall that ${\mathbf{x}}^{\sigma}$ is subject to $p_{\text{data}}({\mathbf{x}})*{\mathcal{N}}(0,\sigma^{2}{\bm{I}})$. Hence, two adjacent discretized samples on the same ODE trajectory can be represented as ${\mathbf{x}}+\sigma_{n}{\mathbf{z}}$ and ${\mathbf{x}}+\sigma_{n+1}{\mathbf{z}}$ where ${\mathbf{x}}\sim p_{\text{data}}$ and ${\mathbf{z}}\sim{\mathcal{N}}(\bm{0},{\bm{I}})$. We can apply the self-consistency condition by minimizing the error between outputs of two adjacent samples. Therefore, the objective function of training CMs is
@@
Consistency Models as Visuomotor Policies.
By analogy to Diffusion Policy [1], CMs can also be adapted to IL settings, serving as robot visuomotor policies. In this case, a CM characterizes the conditional distribution $p({\mathbf{a}}_{t}|{\mathbf{o}}_{t})$ of robot action sequences ${\mathbf{a}}_{t}$ given observation ${\mathbf{o}}_{t}$ from demonstrations. At $t$-th step of closed-loop control, a consistency policy $\bm{\pi}$ predicts an action sequence by evaluating ${\mathbf{a}}_{t}=\bm{\pi}_{\bm{\theta}}(\hat{{\mathbf{a}}_{t}}^{\sigma_{\text{max}}},{\sigma_{\text{max}}}\mid{\mathbf{o}}_{t})$ once. The key distinguishing feature of consistency policies, compared to diffusion policies, is their ability to generate an action sequence through only one-step inference without iterative sampling, thereby being faster while maintaining modeling quality. This makes consistency policies computationally efficient during execution, enabling them to complete dynamic tasks besides quasi-static ones.
@@ 4 Bimanual Keypose-Conditioned Consistency Policy
Problem Formulation.
We assume a set of demonstrated trajectories ${\mathcal{D}}=\{\tau^{i}\}_{i=0}^{D-1}$ collected via teleoperation [33], where $\tau^{i}$ is the $i$-th sequence of observation-action pairs, i.e., $\tau^{i}=\{({\bm{o}}_{t}^{i},{\bm{a}}_{t}^{i})\}_{t=0}^{T_{i}-1}$. The observation ${\bm{o}}_{t}^{i}=({\bm{x}}_{t}^{i},{\bm{q}}_{t}^{i})$ includes RGB images ${\bm{x}}_{t}^{i}$ and proprioceptive information ${\bm{q}}_{t}^{i}$ (i.e., joint positions and gripper width), while the action ${\bm{a}}_{t}^{i}$ is the expected proprioceptive status. Note that we omit the superscript $i$ for simplicity if the context is clearly within one trajectory. In any trajectory $\tau^{i}$, we assume a set of keyposes ${\mathcal{K}}^{i}=\{k_{0}^{i},\dots,k^{i}_{j},\dots,k_{M_{i}}^{i}\}$ exists, which can be represented in either SE(3) or the configuration space. We choose the latter one in this work (i.e., joint positions of bimanual arms) to align with the representation of states and actions. With ${\mathcal{D}}$ and ${\mathcal{K}}$, we aim to learn a keypose predictor ${\bm{g}}_{\bm{w}}({\bm{o}},k)$ for high-level guidance and a trajectory generator $\bm{\pi}_{\bm{\theta}}({\mathbf{a}}|{\mathbf{o}},k)$ producing action sequences ${\mathbf{a}}$ for task fulfillment conditioned on historical observations ${\mathbf{o}}$ and target keypose $k$. We introduce design details of the keypose predictor and the trajectory generator below.
@@ 4.1 Keypose Predictor
Bimanual Keypose Identification.
We summarize and generalize the heuristics in previous studies [9, 31, 23] into three categories to decide whether a robotic arm is at a keypose: (1) changes in robot-object contact mode; (2) motion stalling and (3) case-specific rules. In more detail, changes in contact mode always come with the opening or closing of robotic grippers, which indicates changes in positional relationships among objects and changes in workspace state. Motion stalling means the arm is performing pre-grasp or pre-release, which are significant for fine-grained manipulation. Case-specific rules are dependent on spatial-temporal relationship between two arms in certain tasks, e.g., the relative distance between left and right grippers and the height of grippers w.r.t. the table-top. Specifically, we scan each demonstration trajectory and check whether (1) the gripper begins to open or close; (2) the velocity of gripper is below a threshold; and (3) the relative distance or height is below a threshold and mark those timesteps that satisfy the conditions as keyposes.
@@
Training Keypose Predictor.
We can extract bimanual keyposes ${\mathcal{K}}^{i}=\{k_{j}^{i}\}_{j=0}^{M_{i}}$ from each demonstration $\tau^{i}=\{({\bm{o}}_{t}^{i},{\bm{a}}_{t}^{i})\}_{t=0}^{T_{i}-1}$. Here, $k_{j}^{i}$ is the $j$-th keypose in the $i$-th trajectory in the form of joint positions such that $k_{j}^{i}={\bm{q}}_{t_{j}}^{i}$, and $t_{j}$ is the corresponding timestep in the trajectory ($t_{0}=0,t_{M_{i}}=T_{i}$). Note that the number of keyposes is much smaller than the number of timesteps (i.e., $M_{i}\ll T_{i}$). Therefore, we enrich the training data for keypose predictor by training it to predict target keyposes at all timesteps, rather than only at the designated keypose steps. This approach enhances the predictor’s capability by leveraging a larger volume of training data. Specifically, the keypose dataset is constructed as
@@ 4.3 Implementation and Training Details
Design choices of $\bm{\pi}_{\bm{\theta}}$ generally follows those of CNN-based diffusion policy proposed in [1]. However, a significant modification we made is that, besides visual and proprioceptive features, the 1D temporal CNN backbone also conditions on the target keypose. We also use the same settings of $H_{o}=2$ and $H_{a}=8$. On the consistency training side, we leverage part of improved techniques proposed in [26] to enhance experimental results. Specifically, the weighting function is $\lambda(\sigma_{n})=1/(\sigma_{n+1}-\sigma_{n})$ while the distance metric is Pseudo-Huber loss $d({\bm{x}},{\bm{y}})=\sqrt{\|{\bm{x}}-{\bm{y}}\|^{2}_{2}+c^{2}}-c$ with $c=0.0064$. Furthermore, we adopt $\mu=0$ and the step-shaped discretization curriculum. Both models are trained by AdamW [20] optimizer, with batch size $B=64$ and learning rate cosine annealing from 1e-4 to zero across all tasks. Each element of inputs to both models are normalized to $[-1,1]$ by element-wise maximum and minimum. Then outputs will be unnormalized back to original ranges. More detailed parameters can be found in Appendix B.

### v2

@@ Abstract
Bimanual manipulation tasks typically involve multiple stages which require efficient interactions between two arms, posing step-wise and stage-wise challenges for imitation learning systems. Specifically, failure and delay of one step will broadcast through time, hinder success and efficiency of each sub-stage task, and thereby overall task performance. Although recent works have made strides in addressing certain challenges, few approaches explicitly consider the multi-stage nature of bimanual tasks while simultaneously emphasizing the importance of inference speed. In this paper, we introduce a novel keypose-conditioned consistency policy tailored for bimanual manipulation. It is a hierarchical imitation learning framework that consists of a high-level keypose predictor and a low-level trajectory generator. The predicted keyposes provide guidance for trajectory generation and also mark the completion of one sub-stage task. The trajectory generator is designed as a consistency model trained from scratch without distillation, which generates action sequences conditioning on current observations and predicted keyposes with fast inference speed. Simulated and real-world experimental results demonstrate that the proposed approach surpasses baseline methods in terms of success rate and operational efficiency. Codes are available at https://github.com/ManUtdMoon/BiKC.
@@ 1 Introduction
Bimanual manipulation are ubiquitous in daily activities and industrial applications such as cloth folding, battery slotting and tool assembling [3, 37, 12, 1]. These tasks typically require intricate coordination between two arms and involve multiple stages, resulting in more complexity than unimanual case  [8, 34]. One intuitive solution is to assign a specific role to each arm based on human-specified rules, with each arm being programmed independently [19, 11, 21]. However, this approach proves to be labor-intensive and hard-coded, as coordination rules often require customization for specific tasks, particularly those involving multiple stages [22]. Recently, imitation learning (IL), a.k.a. learning from demonstration (LfD), has been widely applied in bimanual manipulation. IL is a straightforward yet effective approach that obtains automated agents in an end-to-end manner by imitating demonstrations, thereby eliminating the need for explicitly and manually designing coordination rules [4, 36, 20].
@@ 2 Related Work
Imitation Learning for Robotic Manipulation. IL enables robots to acquire skills by learning directly from experts’ demonstrations. The intuitive process makes it widely used in robotic manipulation. IL in manipulation can be categorized into non-parametric approaches (e.g., Dynamic Movement Primitives [14]) and parametric methods promising for high-dimensional tasks [36, 32]. Behavioral cloning (BC) [5] is a classic parametric IL algorithm that acts similarly as supervised learning, directly fitting the mapping from observations (inputs) to actions (labels). Despite its simplicity and efficacy, BC-based approaches are susceptible to compounding errors [37] and multi-modality in demonstrations [2], which is worse when it comes to bimanual cases. Recently, the integration of generative modeling techniques into IL has brought success [3, 37, 2, 28]. Predicting action sequences in these works mitigates compounding errors by shortening the effective horizon [37, 2, 28]. Some works [37, 28] address the multi-modality challenge by leveraging Variational Auto-Encoder (VAE) [18] to focus on the predominant behavior style. Conversely, other works capture multi-modal action distributions by modeling the policy as a diffusion model [2, 13, 31]. However, diffusion models inherently incur significant inference latency due to iterative denoising, which presents a critical concern especially in real-world deployment. In this work, we focus on a variant of diffusion models, Consistency Models (CMs) [30, 29], to formulate robot policies. CMs can capture multi-modal demonstrations like DP [30] while generating action sequences with only one-step inference, i.e., fast inference speed. Some concurrent works [24, 27] also apply CM to robotic manipulation tasks with the aim of accelerating run-time inference. However, these approaches distill CMs from well-trained diffusion models (i.e., consistency distillation (CD) [30]) and focus only on unimanual tasks. Here, CD necessitates additional efforts to learn a diffusion model first and depends on the quality of the learned diffusion model. To this end, we adopt consistency training (CT) [30, 29] in this paper. Through comprehensive experiments, we demonstrate that CMs trained from scratch can achieve high-performance deployment and real-time reaction in bimanual tasks.
@@
Hierarchical IL for Multi-stage Tasks. Hierarchical IL (HIL) is proposed to enhance the performance for multi-stage tasks, by decomposing a complex task into smaller and more manageable ones. One line of studies pre-define a set of manipulation primitives, such as picking, placing, moving and stabilizing. Then a high-level policy is learned to sequentially choose primitives to be executed [25]. In bimanual cases, this intuition is implemented by assigning a fixed role for each arm, after which each arm learns an independent policy for its designated role [11]. In this way, the bimanual system is decomposed into two unimanual cases. However, such a method is constrained to specific types of tasks. Another line of work segments the demonstration into sub-stage trajectories. For instance, [38] employs a bottom-up clustering on observations to accomplish unsupervised segmentation, while [26] establishes heuristic rules based on robot kinematics. One technique to segment trajectories is extracting keyposes. Keypose typically refers to the next-best end-effector (EE) pose or joint positions where the arm reaches a specific state, e.g., EE velocity approaching zero or the gripper beginning to open or close. With identified keyposes, IL on a multi-stage task comes down to modeling both the high-level keypose transition and the low-level sub-trajectories connecting two-keypose pairs. Built on this idea, several works have made achievements in unimanual manipulation [10, 35, 26]. However, applying keypose-based approaches in bimanual tasks poses a challenge in coordinating two arms. In this study, we address this issue by adopting a merging method. Furthermore, we condition the learning of the trajectory generator on target keyposes, to improve per-stage reliability and enhance overall success rates.
@@ 3 Preliminaries: Consistency Models as Visuomotor Policy
Consistency Models in Generative Modeling. CMs are proposed to remove the iterative sampling process in DDPMs to accelerate generation speed [30] as another score-based generative modeling method. The core of CMs is the existence of an ordinary differential equation (ODE) shares the same per-step distribution $p_{\sigma}({\mathbf{x}})$ of samples ${\mathbf{x}}^{\sigma}$ with diffusion SDE [31]:
@@
where ${\bm{f}}(\cdot,\cdot)$ is the consistency function mapping diffused samples to origin.
Training and Sampling of Consistency Models. Training CMs is to approximate the consistency function ${\bm{f}}$ with a parameterized one ${\bm{f}}_{\bm{\theta}}$ (typically neural networks) by learning to enforce eq. 2. Recall that ${\mathbf{x}}^{\sigma}$ is subject to $p_{\text{data}}({\mathbf{x}})*{\mathcal{N}}(0,\sigma^{2}{\bm{I}})$. Hence, two adjacent discretized samples on the same ODE trajectory can be represented as ${\mathbf{x}}+\sigma_{n}{\mathbf{z}}$ and ${\mathbf{x}}+\sigma_{n+1}{\mathbf{z}}$ where ${\mathbf{x}}\sim p_{\text{data}}$ and ${\mathbf{z}}\sim{\mathcal{N}}(\bm{0},{\bm{I}})$. We can apply the self-consistency condition by minimizing the error between outputs of two adjacent samples. Then, the training objective of CMs is
@@
Consistency Models as Visuomotor Policies. By analogy to Diffusion Policy [2], CMs can also be adapted to IL settings, serving as robot visuomotor policies. In this case, a CM characterizes the conditional distribution $p({\mathbf{a}}_{t}|{\mathbf{o}}_{t})$ of robot action sequences ${\mathbf{a}}_{t}$ given observation ${\mathbf{o}}_{t}$ from demonstrations. At $t$-th step of closed-loop control, a consistency policy $\bm{\pi}$ predicts an action sequence by evaluating ${\mathbf{a}}_{t}=\bm{\pi}_{\bm{\theta}}(\hat{{\mathbf{a}}_{t}}^{\sigma_{\text{max}}},{\sigma_{\text{max}}}\mid{\mathbf{o}}_{t})$ once. The key distinguishing feature of consistency policies, compared to diffusion policies, is their ability to generate an action sequence through only one-step inference without iterative sampling, thereby being faster while maintaining modeling quality. This makes consistency policies computationally efficient during execution, enabling them to complete dynamic tasks besides quasi-static ones.
@@ 4 Bimanual Keypose-Conditioned Consistency Policy
Problem Formulation. We assume a set of demonstrated trajectories ${\mathcal{D}}=\{\tau^{i}\}_{i=0}^{D-1}$ collected via teleoperation [37], where $\tau^{i}$ is the $i$-th sequence of observation-action pairs, i.e., $\tau^{i}=\{({\bm{o}}_{t}^{i},{\bm{a}}_{t}^{i})\}_{t=0}^{T_{i}-1}$. The observation ${\bm{o}}_{t}^{i}=({\bm{x}}_{t}^{i},{\bm{q}}_{t}^{i})$ includes RGB images ${\bm{x}}_{t}^{i}$ and proprioceptive information ${\bm{q}}_{t}^{i}$ (i.e., joint positions and gripper width), while the action ${\bm{a}}_{t}^{i}$ is the expected proprioceptive status. Note that we omit the superscript $i$ for simplicity if the context is clearly within one trajectory. In any trajectory $\tau^{i}$, we assume a set of keyposes ${\mathcal{K}}^{i}=\{k_{0}^{i},\dots,k^{i}_{j},\dots,k_{M_{i}}^{i}\}$ exists, which can be represented in either SE(3) or the configuration space. We choose the latter one in this work (i.e., joint positions of bimanual arms) to align with the representation of states and actions. With ${\mathcal{D}}$ and ${\mathcal{K}}$, we aim to learn a keypose predictor ${\bm{g}}_{\bm{w}}({\bm{o}},k)$ for high-level guidance and a trajectory generator $\bm{\pi}_{\bm{\theta}}({\mathbf{a}}|{\mathbf{o}},k)$ producing action sequences ${\mathbf{a}}$ for task fulfillment conditioned on historical observations ${\mathbf{o}}$ and target keypose $k$. We introduce design details of the keypose predictor and the trajectory generator below.
@@ 4.1 Keypose Predictor
Bimanual Keypose Identification. We summarize and generalize the heuristics in previous studies [10, 35, 26] into three categories to decide whether a robotic arm is at a keypose: (1) changes in robot-object contact mode; (2) motion stalling and (3) case-specific rules. In more detail, changes in contact mode always come with the opening or closing of robotic grippers, which indicates changes in positional relationships among objects and changes in workspace state. Motion stalling means the arm is performing pre-grasp or pre-release, which are significant for fine-grained manipulation. Case-specific rules are dependent on spatial-temporal relationship between two arms in certain tasks, e.g., the relative distance between left and right grippers and the height of grippers w.r.t. the table-top. Specifically, we scan each demonstration trajectory and check whether (1) the gripper begins to open or close; (2) the velocity of gripper is below a threshold; and (3) the relative distance or height is below a threshold and mark those timesteps that satisfy the conditions as keyposes.
@@
Training Keypose Predictor. We can extract bimanual keyposes ${\mathcal{K}}^{i}=\{k_{j}^{i}\}_{j=0}^{M_{i}}$ from each demonstration $\tau^{i}=\{({\bm{o}}_{t}^{i},{\bm{a}}_{t}^{i})\}_{t=0}^{T_{i}-1}$. Here, $k_{j}^{i}$ is the $j$-th keypose in the $i$-th trajectory in the form of joint positions such that $k_{j}^{i}={\bm{q}}_{t_{j}}^{i}$, and $t_{j}$ is the corresponding timestep in the trajectory ($t_{0}=0,t_{M_{i}}=T_{i}$). Note that the number of keyposes is much smaller than the number of timesteps (i.e., $M_{i}\ll T_{i}$). Therefore, we enrich the training data for keypose predictor by training it to predict target keyposes at all timesteps, rather than only at the designated keypose steps. This approach enhances the predictor’s capability by leveraging a larger volume of training data. Specifically, the keypose dataset is constructed as
@@ 4.3 Implementation and Training Details
Design choices of $\bm{\pi}_{\bm{\theta}}$ generally follows those of CNN-based diffusion policy proposed in [2]. However, a significant modification we made is that the 1D temporal CNN backbone conditions on target keyposes besides visual and proprioceptive features. We also use the same settings of $H_{o}=2$ and $H_{a}=8$. On CM training side, we leverage part of improved techniques proposed in [29] to enhance experimental results. Specifically, the weighting function is $\lambda(\sigma_{n})=1/(\sigma_{n+1}-\sigma_{n})$ while the distance metric is Pseudo-Huber loss $d({\bm{x}},{\bm{y}})=\sqrt{\|{\bm{x}}-{\bm{y}}\|^{2}_{2}+c^{2}}-c$ with $c=0.0064$. Furthermore, we adopt $\mu=0$ and the step-shaped discretization curriculum. Both models are trained by AdamW [23] optimizer, with batch size $B=64$ and learning rate cosine annealing from 1e-4 to zero across all tasks. Each element of inputs to both models are linearly normalized to $[-1,1]$. Then outputs will be unnormalized back to original ranges. More detailed parameters can be found in Appendix B.